A wave on a string is described by the relation y =0.15 sin(200πt -0.045x), where t is

measured in seconds and x in meters.

(a) What is the time required for a single wavelength to pass a fixed point by the wave?

(b) What is the maximum speed of a point on the string?


1
Expert's answer
2021-01-22T15:04:42-0500

(a) "T=1/\nu=2\pi/\omega=2\cdot\pi/(200\pi)=0.01(s)"


(b) "v=dy/dt=0.15\cos(200\pi t-0.045x)\cdot200\pi=94.25\cos(200\pi t-0.045x)"


"v_{max}=94.25(m\/s)"